OpenStudy (anonymous):

How do I simplify -5ab(2a^2 B^3)

OpenStudy (just_one_last_goodbye):

hmm let me get on one ok

OpenStudy (anonymous):

Yeth

OpenStudy (anonymous):

Just realized, youre a human calculator

OpenStudy (anonymous):

help me out

OpenStudy (anonymous):

you must be a genius

OpenStudy (anonymous):

-10a^3b^4

OpenStudy (anonymous):

ez pz

OpenStudy (anonymous):

can you explain to me how you got that answer pls?

OpenStudy (anonymous):

I have my notebook rdy

OpenStudy (anonymous):

-5(2) = -10 a(a^2)= a^3 b(b^3) = b^4

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):



OpenStudy (anonymous):